OpenStudy (anonymous):

What is the slope of the line that passes through (2, 5) and (−1, 5)?

OpenStudy (anonymous):

\[slope = \frac{rise}{run}\]

OpenStudy (akashdeepdeb):

If you are given 2 points \[(x_1,y_1) ~~and~~~ (x_2,y_2)\] Their slope is given by \[\frac{Change~~in~~ Y}{Change~~in~~X}\] \[=\frac{y_2-y_1}{x_2-x_1}\] Understood? :)

OpenStudy (anonymous):

thank you (: